OpenStudy (anonymous):

If a=4, b=2band c= -3, find 6a-b

OpenStudy (anonymous):

I think the answer is 22 but im not sure

OpenStudy (anonymous):

6*4=24-2=22

OpenStudy (anonymous):

thanks i wasnt sure